OpenStudy (anonymous):

Calculate molality of 1litre solution of 93% H2So4 by volume. The density of solution is 1.84 g/ml.

OpenStudy (australopithecus):

First the formula for Molality is: molality = moles of solute/kg of solvent So you need to find (1) moles of Solute, and (2) and calculate kg of solvent 1. Calculate moles of H2SO4 a) Determine how many moles of H2SO4 are in 1L of solution (since it says in the quesstion we have 1L of solution) (we are taking H2SO4 as the solute and water as solvent) we know it is a 93% volume solution Formula for % volume solution is: % solution = (Volume solute/ Volume of Solution)*100 (NOTE: volume of solute (H2SO4) and volume of solution (1L) must have the same units, can be L or mL but they have to be the same otherwise the units wont cancel and you wll be off by some factor of 10, or maybe something different if you don't use metric) You know the percent solution and you know the volume of solution so just solve for volume of solute You need to convert the volume of H2SO4 to mass, so look up the density of H2SO4 (I did it for you): Density of H2SO4 = 1830.5g/L To solve for mass given any volume of H2SO4 using the density set up the rato: \[\frac{1830.5g}{1 L} = \frac{x \ g}{y \ L}\] You need to solve for x in this case, as you know y it is the volume you solved for using the % volume formula. You should now know your mass of H2SO4 in 1L of solution in grams. b) Calculate molar mass of H2SO4 using periodic table You can look this up if you dont know how to do it (Just search on youtube calculating molar mass of a molecular using the periodic table) Using the Mass of H2SO4 in the solution you just calcualted use the following formula to solve for the total moles Moles = grams/Molecular mass you know grams and you know the molecular mass of H2SO4 so just solve for moles of H2SO4 write this down and circle it you need it 2. Calculate Mass of Solvent (water is your solvent) Realize this: Total Mass of Solution = Mass of Solute + Mass of Solvent We know the mass of solute (we calculated that) We want to find the mass of Solvent (remember it has to be in kg) We can find the total mass of solution using the density of solution given in the question: Density of solution = 1.84 g/ml We just do what we did earlier set up an equality, we \[\frac{1.84g}{1 mL} = \frac{x \ g}{y \ mL}\] you know y it is 1L (the total volume of solution), you just need to solve for x. Note the units are in mL not in L. Make sure you convert units. If you are confused about unit conversion just look it up on youtube, or ask for help. Now you know the total mass of solution and can solve for mass of solvent using: Total Mass of Solution = Mass of Solute + Mass of Solvent 3. Sub in the moles of solute (H2SO4) and Mass of Solvent in kg into the formula to solve for molality (it uses the unit lower case m, for example 1m = 1 molality): molality = moles of solute/kg of solvent Hope this helps

OpenStudy (anonymous):

thnx a lot:)